Solve by completing the square. 4h2 + 16h - 20 = 0 h= (Simplify your answer. Type an exact answer, using
Answer #1

Given, 44h+164- 20 = 0 +(4+44-5) = 0 Now, h44h-5=0 H-1 + 5h - 5 = 0 (4-1)+5(1-1)0 (h+5)(n-1) = 0 ... 1+5 = 0 ; H-1 - 0 he-5 ;